Mr. Rahami invested $15,000, a part at 8% and the rest at 11%. If his yearly income from these investments is $1335, how muchdid he invest at each rate ( assuming simple interest)?
Answer by ikdeep(226)

let money invested in one part = $ x
and let money invested on other part = $y
according to the first condition i.e. Mr. Rahami invested $15,000,,we get the equation....
x + y = $15,000.........(1)
according to the second condition i.e. a part at 8% and the rest at 11%. If his yearly income from these investments is $1335,,we get the equation....
(8% of x) + (11% of y) = $1335
or 0.08x + 0.11y = 1335............(2)
now solve equation (1) and (2) ny elimination method ,,and you will get the answer....
Still if you are unable to solve further, contact me again.
hope this will help you
Please feel free to revert back for any further queries.